OpenStudy (anonymous):

May someone tell me I worked out this problem correctly f(x) = 7x -2x ^2 where f(x-2) 1) f(x-2) =7(x-2)-2(x-2)^2 2) f(x-2) = 7x-14-2(x^2-4x+4) 3) f(x-2) = 7x - 14 -2x^2+8x-8 4) -x-22-2x^2 For some reason my solution doesn't look right to me...

OpenStudy (anonymous):

F(x)=7x-2x^2 F(x-2)=7(x-2)-2(x-2)^2 F(x)=7x-14-2(x^2+4 -4x) F(x)=7x-14-2x^2-8+8x It's third option.

OpenStudy (amistre64):

it seems right so far

OpenStudy (anonymous):

it needs to be simplified though, which is where I got the -x-22-2x^2

OpenStudy (amistre64):

7x - 14 -2x^2+8x-8 -2x^2 +x(8+7) -(14+8)

OpenStudy (anonymous):

Sorry I failed...thought they're options.. -2x^2+15x-22

OpenStudy (amistre64):

when you combined terms you did a nono :) they are spose to be LIKE terms

OpenStudy (anonymous):

@Amistre: -2x^2+15x-22 can't this be the final term?

OpenStudy (amistre64):

it can be, but it depends on how anal the grader is

OpenStudy (amistre64):

the negation of it would seem more proper cause they dislike a negative leading coeff

OpenStudy (anonymous):

the grader isn't too anal, just as long as I show my work and prove how I got to that solution. That is the originial solution I got, but then I starte second guessing myself

OpenStudy (anonymous):

@amistre: how 'anal'?..